### v1

# Literature Review: Cyber Security Monitoring in Maritime

Abstract
In recent years, many cyber incidents have happened in the maritime sector, targeting the information technology (IT) and operational technology (OT) infrastructure. Although several systematization-of-knowledge papers have been published in the maritime field, none of the previous studies has focused on cyber security monitoring, which aims at timely detection of cyber attacks with automated methods.
The current article addresses this research gap and surveys the methods, algorithms, tools and architectures used for cyber security monitoring in the maritime sector. For the survey, a systematic literature review of cyber security monitoring studies is conducted in this article, following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) protocol.
The first contribution of this article is the bibliometric analysis of related literature and the identification of the main research themes in previous works. For that purpose, our article presents a taxonomy for existing studies which highlights the main properties of maritime cyber security monitoring research. The second contribution of this article is an in-depth analysis of previous works and the identification of research gaps and limitations in existing literature. Based on our findings, we outline future research directions for cyber security monitoring in the maritime field.
Index Terms:
cyber security, cyber security monitoring, intrusion detection, maritime, monitoring, literature review
I Introduction
Modern societies are heavily depending on shipping and port industries, and maritime services and infrastructures are often regarded as critically important [1]. During the recent decades, maritime systems have become increasingly digitalised and interconnected [2, 3, 4]. For example, a modern vessel network relies heavily on network technologies such as Ethernet to connect the components of the ship’s navigational system. As another example, the ship’s Cyber-Physical System (CPS) is similarly dependent on solutions and technologies found in traditional IT systems.
Reliance on IT technologies has introduced similar cyber security issues into the maritime domain that can be found in traditional IT systems – like organisational computer networks, maritime systems are susceptible to cyber attacks that can disrupt normal operations for longer periods, thus inflicting significant financial damage and potentially leading to catastrophic consequences. For example, in 2017, the Maersk logistics company network was attacked by NotPetya malware which brought many port terminals of the company to a halt [5]. Although the company managed to resume normal operations swiftly, it was estimated that the financial losses inflicted by the malware could have reached as high as 300 million US dollars [6].
Unfortunately, the above incident is not a rare event, but several similar cyber attacks have been conducted against maritime infrastructures. For example, in a recent paper Afenyo and Caesar describe 12 similar security incidents in 15 large ports [7], and according to their report, in 2017 large ports experienced 12 cyber attacks per day. Also, the average ransom paid due to ransomware attacks against maritime systems is 3.1 million US dollars [7]. Additionally, cyber attacks directly affecting the control of ships or monitoring of ship traffic have been recently reported. For example, in 2016 a cyber attack misdirected two navy vessels in the Persian Gulf [8]. In a case from 2017, cybercriminals gained access to the navigation systems of a container vessel owned by a German company, which had a capacity of 8250 TEUs [9]. Moreover, an Italian base station of Automatic Identification System (AIS) experienced a ship spoofing incident near Elba Island, where thousands of fake ships suddenly appeared and affected vastly the accurate monitoring of the maritime traffic in the vicinity [10].
In response to the escalating cyber threats in the maritime sector, leading maritime organisations like the International Maritime Organisation (IMO) and the Baltic and International Maritime Council (BIMCO) have been proactive in developing guidelines and recommendations to bolster maritime cyber security. The IMO, in particular, has integrated cyber security into its safety management systems under the ISM Code. This integration, effective from January 2021, mandates ship owners and managers to assess cyber security risks and develop necessary countermeasures to address these risks as part of their safety management systems [11]. BIMCO, the world’s largest international shipping association, has also released several guidelines on cyber security, including the ’Guidelines on Cyber Security Onboard Ships’ [12]. These guidelines are designed in collaboration with other industry associations and bring forward a comprehensive framework for cyber security risks management in maritime operations. The framework covers various aspects of cyber security, from identifying threats and vulnerabilities to implementing effective risk management processes.
Despite the recent incidents and the increasing volume of recommendations and guidelines by leading maritime organisations, Kechagias et al. [2] pointed out that cyber security remains a relatively new concept for the maritime industry, with many companies having low cyber security awareness and immature cyber security risk management culture. Furthermore, Afenyo and Caesar state in their paper [7] that current study programs of maritime educational institutions offer insufficient cyber security knowledge to their students and trainees. These factors have led to low adoption of cyber attack detection and mitigation solutions in the maritime domain and to insufficient amount of academic research conducted in this field.
Cyber attack detection and mitigation is further complicated by unique features of maritime systems (see [13, 14, 15] for a more detailed discussion). Maritime systems usually consist from a number of different IT and OT networks that connect a wide variety of maritime devices which communicate over specialised protocols such as NMEA. Therefore, the detection of attacks conducted in such networks requires detailed analysis of these maritime-specific protocols by dedicated solutions. Furthermore, ships have a limited network connectivity with the shore, which leads to the need for autonomous attack detection functionality on vessels. Also, space limitations and other physical constraints do not often allow to operate significant computational resources (e.g., server clusters) on board of ships, introducing the need for lightweight and resource-efficient cyber security solutions.
One of the key approaches for countering cyber attacks is cyber security monitoring (or security monitoring for the sake of brevity), which aims at detecting cyber attacks with automated methods in a timely fashion, allowing to mitigate the attacks in a manual, semi-automated, or fully automated way. For example, network and host-based Intrusion Detection Systems (IDSs) are widely acknowledged security monitoring technologies for real-time surveillance of computer networks and individual hosts to detect cyber attacks. Also, most security-aware organisations operate a Security Operations Centre (SOC) for collecting data from security monitoring tools to a central Security Information and Event Management (SIEM) system where it is processed and then presented to human security analysts for further action [16]. In addition to existing industrial solutions, many experimental approaches like Machine Learning (ML) based algorithms have been proposed for cyber security monitoring in recent academic literature (see a domain overview paper by Apruzzese et al. [17] for more details).
Although security monitoring technologies have been widely adopted by traditional (i.e., non-maritime) organisations and much academic research has been conducted in this field, security monitoring remains understudied in the maritime domain. For example, detailed cyber security guidelines by BIMCO [12] touch security monitoring only briefly, mentioning the use of network IDS and malware detection for security monitoring purposes, but failing to provide more detailed implementation recommendations to build maritime SOCs and other maritime security monitoring solutions. As another example, cyber risk management guidelines by IMO [18] do not address cyber security monitoring at all. As pointed out in [13], the generic nature of existing guidelines complicates the creation of maritime security monitoring systems, and dedicated research in this domain is needed. Also, as discussed in Section IV of this paper, maritime security monitoring is a relatively new research field, with most academic papers being published in the last 5-6 years. Due to the novelty of this research field, no domain overview studies have been published in maritime security monitoring that would analyse existing academic literature and identify the main research topics and open challenges.
This article fills the aforementioned research gap and specifically targets the research on security monitoring in the maritime domain. The purpose of the current study is to provide a Systematic Literature Review (SLR) of research papers published in the field. Note that our study focuses on peer-reviewed academic papers, excluding grey literature from consideration. A similar approach has been used in recent SLRs that address maritime cyber security, risk, and safety [3, 19]. Using the same approach allows us to evaluate existing literature from an academic perspective and to identify research themes and gaps in analysed papers.
Apart from considering peer-reviewed publications only, we did not impose any additional limitations to the literature selection process. As a result, analysed papers cover a wide variety of maritime cyber security monitoring topics, including the design of a maritime SOC, the creation of testbed environments for maritime security monitoring experiments, and various security monitoring methods (e.g., ML-based approaches) for navigational systems, vessel CPSs, maritime IoT networks, etc.
Our study begins with a discussion of selection criteria for relevant research that is followed by bibliometric analysis. We continue with creating a taxonomy for maritime security monitoring research and describe the trends and common themes in this research. Finally, our study provides a thorough analysis of existing research papers and identifies open issues, outlining future research directions for addressing these research gaps.
The research gaps we have found are the following. First, existing research has not utilised publicly available maritime-specific datasets. Also, many existing papers are suffering from several evaluation issues, for example, failing to employ standard performance metrics, to measure the computational cost of the proposed methods, and to release the experiment code publicly. These deficiencies are in conflict with the best practices of the security monitoring research, decreasing the reproducibility of the experiments and complicating the assessment of the research results [17, 20]. Furthermore, we discovered a number of unexplored research topics during literature analysis. For example, existing studies on maritime cyber security monitoring have not addressed advanced ML issues like concept drift and learning models for resource-constrained environments. As another example, human aspects of maritime security monitoring have received little attention in the existing literature. For a detailed discussion of identified research gaps, please see Section VI.
The remainder of this article is organised as follows. Section II describes similar domain overview and SLR papers, Section III formulates the research questions of our study and discusses the selection criteria for research covered by our study, and Section IV contains the bibliometric analysis of the research papers selected for the study. Section V presents the taxonomy of relevant research and provides an in-depth analysis of selected research papers. Section VI discusses open issues and research gaps in the existing research literature, and Section VII concludes the article.
II Related work
The purpose of this section is to provide an overview of similar SLR papers in the maritime domain. At the end of this section, we will also mention some systematisation-of-knowledge papers on intrusion and anomaly detection in non-maritime environments. The maritime SLR papers discussed in this section are summarised in Table I. As can be seen from Table I, existing SLRs have mainly focused on the analysis of past cyber incidents and ways for mitigating cyber threats in the future. For that purpose, the authors have proposed various measures, most notably by changing existing regulations and policies, and by improving the cyber security training and cyber awareness of personnel. Only one SLR paper [3] covers some security monitoring studies, but their treatment has remained very brief without a deeper analysis. Although the papers from Table I do not focus on maritime security monitoring, we will discuss these papers for the sake of the completeness of our study, and to illustrate that the treatment of maritime security monitoring has remained too brief in existing SLRs.
TABLE I: Maritime Cyber Security SLR Papers
Paper Year Topics Analysis of Security Monitoring
Research Literature
[21] 2023 impact assessment for cyber threats, categorisation and mitigation of cyber risks missing
[22] 2023 cyber risk management, analysis of regulations and policies missing
[7] 2023 analysis of cyber threats, regulations and policies missing
[19] 2023 risk management, reliability and safety of autonomous ships missing
[2] 2022 ensuring cyber security during digitalisation of a marine company missing
[5] 2022 analysis of cyber threats, regulations and policies missing
[3] 2022 maritime cyber security and risk management brief treatment of some studies
without a deeper analysis
[4] 2022 analysis of past cyber incidents and potential cyber attacks against maritime systems missing
[10] 2020 analysis of attacks against navigational systems and mitigation measures missing
Yu et al. [21] conducted an SLR on cyber security in the maritime domain, describing common trends in related literature. As commonly occurring research themes, the authors identified the impact assessment for maritime cyber threats, recognition of maritime cyber threats (e.g., through vulnerability management tools or through interviews with human personnel), and categorisation and mitigation of maritime cyber risks. The study also presented recommendations for risk mitigation, and discussed gaps in existing literature and avenues for further research. As a supplementary material to their paper, the authors published a list of 21 maritime cyber attacks that have received media coverage and public attention, providing a short description of each cyber attack. Although the study briefly mentioned couple of security monitoring approaches for cyber risk mitigation (e.g., the use of ML methods for AIS anomaly detection), these approaches were not elaborated further.
Androjna et al. [10] investigated the prospect of cyber threats and confirmed incidents within the maritime industry, aiming to identify the emerging cyber trends and challenges impacting safe navigation and marine shipping. The authors focused especially on the vulnerabilities of the Global Positioning System (GPS) as one of the core components for safe navigation. An SLR was conducted supplemented by an analysis of the aforementioned specific spoofing event that occurred on Elba Island in 2019. The authors summarised that a change in thinking is important to drive more processes towards cyber security as well as the need for manufacturers to enhance the cyber security of their designs. Encouraged by ongoing national defence research on alternative positioning, navigation and timing systems, as well as by IMO’s e-Navigation concept, the authors also proposed the use of systems that offer different positioning, navigation, and timing capabilities aboard ships, serving as to facilitate GPS-inclusive navigation. Furthermore, the authors recommended utilising various satellite navigation constellations, both public and private, alongside earthbound sub-systems of an enhanced LOng-RAnge Navigation (eLORAN) system as well as advanced ship aiding systems at ports. This approach aims to facilitate the maritime domain with the required improved defense against cyber threats in the near future.
Kechagias et al. [2] reviewed the aspects that should be considered during the digitalisation of a marine company. The review was focused on cyber security within the maritime domain, presenting an observation of the current state of the topic. The authors described how real-world organisations can manage and develop overall consciousness and intelligence in safeguarding shipping from malicious actors. The significance and main attributes of cyber security in the maritime domain were introduced and analysed. The authors outlined systemic methodologies by pointing out improved procedures and policies. Initially the plan-do-check-act (PDCA) method was introduced for emphasising continual enhancement. Another method then emphasised the effective management of the three primary aspects of security i.e. procedures, the human factor, and technology. Additionally, to augment comprehension of the topic, an audit survey of vessel cyber security was conducted.
Schinas and Metzgera [22] discussed maritime cyber risks, the IMO instruments, the trend of digitalisation, and insurance-related literature. The authors revisited the concept of seaworthiness and examined it from the perspective of cyber risks with a special focus on maritime autonomous surface ships, figuring out the importance of the regulations’ update. As a response to the identified policy gap, the authors introduced the term “cyber-seaworthiness”.
Afenyo and Caesar [7] focused on maritime cyber security threats in their overview paper, analysing relevant literature in the field. According to the authors, existing papers do not provide enough data on past cyber security incidents and this complicates research on modelling these attacks. In addition, the authors stated that the research on maritime cyber attack data-sharing methods has received little attention so far. Furthermore, it was stated that existing studies have not focused on assessing the financial impact of cyber attacks. The authors also noted that the current educational system does not prepare maritime professionals well enough to handle cyber threats. Finally, the authors pointed out that existing governmental and international policies do not regulate maritime cyber security well enough. For example, regulations should facilitate timely reporting of cyber security incidents and cooperation between maritime companies to increase knowledge sharing on the nature and mitigation of cyber attacks.
In their paper, Bronk and deWitte [5] described cyber issues of maritime vessels and ports, providing examples of critical shipboard navigational systems and vulnerable software solutions used in ports. The authors also reviewed some maritime related cyber security policies and guidelines. For example, the US Coast Guard Cyber Strategy emphasises the importance of risk management which relies on promoting cyber risk awareness and reduction of vulnerabilities. Also, IMO’s cyber risk management guidance describes eight areas which may contain vulnerable systems such as bridge systems, cargo handling systems, public networks designed for serving passengers, etc. The authors concluded by identifying research on protecting the IT systems of ships and ports as an important future research direction.
Bolbot et al. [3] conducted an SLR for maritime cyber security. The review included a bibliometric analysis that explored the countries, authors, academic venues and prevalent topics of the publications within the scope. The research studies were categorised and reviewed under topics ranging from various technical topics (e.g., cyber risk management, design of countermeasures, penetration testing) to many interdisciplinary topics such as maritime law, training development, and cyber incident analysis. Cyber risk management constituted a significant portion of the reviewed studies. Although some security monitoring papers were discussed in the study, their treatment remained very brief, since the focus of the article was lying elsewhere.
Another SLR by Chaal et al. [19] analysed the research literature on the risk, reliability, and safety of autonomous ships. Whereas cyber security was not the primary focus of this SLR, it was identified as an important part of the ship’s safety. Also, the study indicated that the cyber security of autonomous vessels has started to receive more attention from researchers recently. However, the study did not include the treatment of cyber security monitoring topics.
Another study by Ben Farah et al. reviewed cyber incidents in the maritime domain and provided a relatively concise bibliographic analysis [4]. The authors provided an extensive elaboration of maritime sub-systems located in vessels and ports, in addition to a review of the potential cyber attacks targeting these systems. Similarly to previous studies described in this section, the authors did not focus on cyber security monitoring issues.
Intrusion and anomaly detection have been broadly addressed in various review studies without focusing on specific systems or network types [23, 24, 25]. Following the introduction of the Internet of Things (IoT) networks, a huge body of literature has researched methods and approaches to run intrusion detection functions within the limits of this network category. Some papers reviewed the relevant literature and discussed the challenges and solutions for IoT environments [26, 27]. Vehicular ad hoc networks (VANETs) and medical systems are also other important system categories that have recently been the subject of various literature review papers in the intrusion detection domain [28, 29].
Our article addresses a gap in the existing SLR literature on the cyber security of the maritime sector. Unlike previous analyses that studied various aspects, our contribution specifically delves into maritime security monitoring systems, an area that has not been adequately covered thus far.
III Methodology
In this section, the systematic review process will be explicated, describing the undertaken methodological approach together with the inclusion and exclusion criteria for relevant research, as well as the overarching review objectives.
III-A Aims and research questions
An SLR is a methodological approach for collecting studies that constitute the research literature in the given field, systematically investigating this research field, answering focused research questions, and identifying the research gaps and research directions in the field [30]. Concretely, this article provides answers to the following research questions:
RQ1: What are the statistical characteristics of the papers that constitute the relevant literature? Analysing the literature through bibliometric methods will provide insights on the publication timeline, the types of venues in which the works were presented, the most prolific authors, and other properties of related work. RQ1 is answered by the discussion in Section IV.
RQ2: What research themes and topics are the papers focusing on, and what is the nature of research presented in the papers? The in-depth analysis of the paper topics will be supported by the paper taxonomy created by us, which is essential for facilitating a meaningful discussion. RQ2 is answered by the analysis in Section V.
RQ3: What are the main shortcomings and limitations of the papers? A thorough examination of the papers will reveal potential research gaps that can be addressed in future studies. RQ3 is answered by the discussion in Section VI.
III-B Survey Method
During the development of the review the PRISMA protocol [31, 32] was followed. According to the protocol, an SLR should have clearly defined objectives or research questions (see Section III-A), eligibility criteria for relevant literature (discussed in Section III-B1), a search strategy for finding relevant studies and selection process (discussed Section III-B2), and data extraction process (discussed in Section III-B3).
III-B1 Eligibility Criteria
Initially, criteria need to be set that guide the search process and help to filter out irrelevant works. In this article, papers written in English and published in peer-reviewed journals and conference/workshop proceedings are included. When searching for relevant papers, we did not set an age limit for publications (i.e., we did not exclude papers published before a specific date). Papers that did not focus on maritime cyber security monitoring were excluded, just like papers focusing on non-IT-related methods from the fields of electrical engineering, physics, and other fields not related to IT. Editorials, opinions, keynotes, abstracts, tutorial summaries, position papers, panel discussions, technical reports and posters were also excluded. The duplicated publications because of different versions were also excluded, with only the most recent version being included. Since this SLR considers peer-reviewed academic papers, grey literature was excluded from this research.
Inclusion and exclusion criteria are summarised below. For a paper to be included, all inclusion criteria described in Section III-B1a must be met. On the other hand, a paper is excluded if it matches any of the exclusion criteria from Section III-B1b.
III-B1a Inclusion Criteria
  • Studies that are peer-reviewed journal and conference papers.
  • Studies in the domain of cyber security in the maritime sector with a focus on monitoring and intrusion detection.
III-B1b Exclusion Criteria
  • Studies that do not explicitly consider cyber security monitoring and intrusion detection.
  • Studies without any connection to maritime.
  • Studies that are not peer-reviewed papers.
  • Grey literature.
  • Studies not written in English.
  • Studies with their full texts not available.
  • Studies that deal with electrical engineering, physics, and other fields not related to IT.
III-B2 Search Strategy
The studies were searched according to the following steps.
III-B2a Search for Related Publications
An initial search was applied to determine the current maritime SLR papers and to identify other significantly related publications. The most notable papers that we found have been described in Section II.
III-B2b Keyword Selection
The selection of keywords heavily influences the results of both automatic and manual searches. During the search phase, several keywords and their combinations were tried on search engines for electronic libraries.
All the search engines that we used support the creation of search queries with Boolean operators (AND, OR and NOT), which helps to narrow the focus of the search and limit the number of hits. Also, the engines allow to apply the search expression on the title, abstract, keywords, metadata, etc. of the papers. The time frame of the search is also definable, and the wildcard characters (such as asterisk (*) and question mark (?)) allow further flexibility during the search.
In the first phase of keyword selection, the keywords related to the maritime, like marine, ship, and navy were selected to define the sector. These keywords combined with the “OR” operator led to papers meant to protect areas against marine infrastructure. Then, the focus was specified —in the same way— by keywords and phrases, like monitor, “intrusion detection”, IDS, “attack detection”, etc. The query made of these keywords brought many publications, but the majority were irrelevant because of the territorial defence mechanisms and different anomalies in the sector. By adding “cyber” to these keywords, the focus was narrowed down sufficiently enough.
III-B2c Automated Search on Bibliographical Sources
During the initial stage of the automated search on bibliographical sources several trials were applied, using different combinations of keywords according to the main objectives of the review, with the aim to construct an optimal keyword set. The final keyword set was then used to construct the search query to be instantiated and executed in various search engines of electronic libraries. The selected electronic libraries include the most important conference proceedings and highest impact journals. The following electronic libraries were used with the same query applied after it was adapted accordingly:
  • ACM digital library (http://portal.acm.org/)
    [[Abstract: monitor*] OR [Abstract: “intrusion detection”] OR [Abstract: “anomaly detection”] OR [Abstract: “attack detection”] OR [Abstract: ids]] AND [Abstract: maritime ship navy navy sea marine ais]
  • IEEE Xplore (http://ieeexplore.ieee.org/Xplore/)
    (“All Metadata”: maritime OR “All Metadata”: ship OR “All Metadata”: navy OR “All Metadata”: sea OR “All Metadata”: marine) AND (“All Metadata”: monitor* OR “All Metadata”: “intrusion detection” OR “All Metadata”: “anomaly detection” OR “All Metadata”: “attack detection” OR “All Metadata”: IDS) AND (“All Metadata”: cyber)
  • Elsevier Scopus (https://www.scopus.com/)
    (TITLE-ABS-KEY ( maritime OR navy OR ship OR marine OR sea OR ais) AND TITLE-ABS-KEY (“intrusion detection” OR monitoring OR “anomaly detection” OR “attack detection” OR ids) AND TITLE-ABS-KEY(cyber))
  • Web of Science (https://www.webofscience.com/)
    cyber (All Fields) AND “intrusion detection” OR monitoring OR “anomaly detection” OR “attack detection” OR ids (All Fields) AND maritime OR navy OR ship OR marine OR sea OR ais (All Fields)
III-B2d Paper Selection Process
The automated search on bibliographical sources resulted in over 1000 papers. After the library search, the four co-authors processed one library each in the form of independent work, and the findings were added to a shared database that helped to indicate the duplicated items. The next step in the selection of relevant papers was the analysis of the title and the abstract. If a paper was irrelevant based on the title and/or the abstract, it was excluded. (In case of doubt, the researcher also read the full text.) If title and abstract were in the scope, the paper became a candidate. Since one more co-author joined to keep the workload balanced, the candidate papers were re-shared among the authors, who reread the full text.
If a paper remained relevant, its main characteristics were noted to support our findings. After the full-text reading of the candidates, we had a set of papers conforming to the inclusion criteria. That constituted the base of the backward snowballing procedure, i.e., additional search by manually checking and analysing the related references. Forward snowballing was also applied to find relevant papers lately published citing any of the selected papers. This additional task ensured the inclusion of every relevant study [33]. The eligibility analysis ended up in 23 relevant papers. At this point, the bibliometric analysis was written, and the methodology for analysis was developed.
III-B3 Data Extraction and Analysis
Following the SLR protocol, at this stage, the papers were studied deeply, summarised, and assessed. The data extracted from each paper were:
  • Publication venue (e.g., name of the journal).
  • Publication time and detailed author information.
  • Research topic of the publication.
  • Introduced monitoring method.
  • Data type processed by the monitoring solution.
  • Data and implementation availability.
  • Performance and computational cost analysis.
  • Validation/evaluation of the presented results.
IV Bibliometric analysis
To answer RQ1 (see Section III-A), this section presents the bibliometric analysis of relevant research.
In light of the information presented, the temporal dissemination of the identified papers is introduced in Figure 1. The annual distribution of published papers does not reveal a discernible trajectory. However, an interesting observation comes up: the number of publications started rising after the major cyber incidents in 2017 within the maritime sector (e.g., the cyber attack against Maersk described in Section I).
Figure 1: Year of publication.
As can be seen from Figure 1, 21 research papers out of the 23 originate from 2018–2023, whereas only two research papers have been published before 2018. In other words, maritime security monitoring is a relatively new research domain which has started to receive more attention during the last 5–6 years.
In accordance with the representation delineated in Table II, the preponderance of contributions assumes the form of conference papers, while about a third of the corpus manifests as journal articles. As for conference papers, it is worthwhile to note that the vast majority of the papers (14 in total) have been published by IEEE, whereas only one paper originates from an ACM conference. When investigating the conference venues more closely, we found that three papers were published at the International Conference on Cyber Situational Awareness, Data Analytics and Assessment, whereas each remaining conference accounted for one paper. As for journal papers, we found that three papers appeared in IEEE Transactions on Intelligent Transportation Systems. Furthermore, Ocean Engineering, Applied Ocean Research, MDPI Journal of Maritime Science and Engineering, MDPI Information, and MDPI Energies published one paper each.
TABLE II: Publications’ type and publisher
Type Total IEEE ACM Elsevier MDPI
Journal article 8 3 - 2 3
Conference paper 15 14 1 - -
Total 23
The 23 research papers were authored by 99 researchers, and we also tried to identify more prolific authors in the field. For that task, we used a frequent itemset mining algorithm to identify all authors who have published at least two papers. Since the purpose of the frequent itemset mining algorithm is to identify associations, that allowed us to detect not just single authors, but also prolific co-author groups. The results are shown in Table III.
TABLE III: Prolific authors
Authors Number of papers
Boudehenn, Boudraa, Cexus 2
Boudvin, Brosset, Kermarrec 2
Jacq 2
Laso 2
As Table III indicates, no author has published more than two papers in the research domain, and a small minority of authors (8 out of 99) have published more than one paper. That indicates that for the majority of authors, maritime security monitoring has not been a persistent focus area with a larger number of peer-reviewed academic publications. In a recent SLR paper by Bolbot et al. [3], a similar phenomenon has been reported – only 17% of authors have published more than one paper in the field of maritime cyber security.
Also, although we identified two co-author groups in Table III, we did not find any group with a substantial number of published papers. Furthermore, as two co-author groups from Table III are coming from the same country, we did not detect any indications of prolific international research cooperation. This illustrates the fact that maritime security monitoring has received modest attention so far, and well-established domestic and international research groups specifically focusing on this particular field have not yet emerged.
V Findings
V-A Taxonomy
To answer RQ2 (see Section III-A), we have created a taxonomy for existing research literature that is presented in Figure 2. Table IV provides detailed information about the selected 23 research papers and how they map to different taxonomy categories. As Figure 2 illustrates, the research papers can be categorised by the environment they target, the type of data which the proposed methods analyse, the type of the proposed methods, and the type of cyber attacks and anomalies that the proposed methods are able to detect. Note that these four categorisations from Figure 2 correspond to the “Targeted Environment”, “Monitored Data”, “Method Type”, and “Detected Attacks and Anomalies” columns in Table IV.
Figure 2: Taxonomy for existing research literature.
TABLE IV: Overview of Research Papers
Paper Year Paper Targeted Monitored Data Method Monitoring Method Detected
Type Environment Type Attacks
and
Anomalies
[34] 2023 M distributed maritime network traffic supervised deep learning IoT attacks
system/network
[35] 2023 M distributed maritime network traffic supervised passive aggressive IoT attacks
system/network machine learning
[36] 2023 M navigational system application layer protocol unsupervised ensemble of spoofing
(NMEA GPS) unsupervised methods
[37] 2023 M distributed maritime network traffic supervised deep learning IoT attacks
system/network
[38] 2023 M distributed maritime network traffic supervised deep learning network
system/network layer attacks
[14] 2022 T distributed maritime various (network traffic, testbed various (IDS signatures, various
system/network NMEA, etc.) SIEM queries, etc.)
[39] 2022 T vessel CPS various (network traffic, testbed various (IDS CPS attacks
fieldbus traffic) signatures, etc.)
[40] 2022 M navigational system application layer protocol rule-based predicate logic rules navigational
(NMEA) anomalies
[15] 2021 M navigational system application layer protocol unsupervised One-Class Support spoofing
(NMEA GPS) Vector Machine
[41] 2021 M vessel CPS OT (ship propulsion system) supervised Particle Filter CPS attacks
[42] 2021 M distributed maritime application layer protocol supervised Light Gradient IoT attacks
system/network (IoT device events) Boosting Machine
[43] 2020 M vessel CPS OT (ship propulsion system) unsupervised Teager-Kaiser operator CPS attacks
for time series analysis
[44] 2020 M vessel CPS OT (ship propulsion system) graph-based graph-based method CPS attacks
[45] 2020 M navigational system application layer protocol rule-based description logic rules navigational
(AIS) anomalies
[46] 2019 M canal CPS OT (canal lock device) supervised deep learning CPS attacks
[47] 2019 A port infrastructure various (network traffic, architecture various (IDS and various
event logs, etc.) antivirus signatures, event
correlation rules, etc.)
[13] 2018 A distributed maritime various (network traffic, architecture various (IDS signatures, various
system/network event logs, etc.) event correlation rules, etc.)
[48] 2018 A distributed maritime various (network traffic, architecture various (agents with various
system/network OT systems, etc.) different functionality)
[49] 2018 M navigational system application layer protocol rule-based distributed rule- spoofing
(AIS) based processing
[50] 2018 A distributed maritime various (ship navigational architecture, Isolation Forest various
system/network data, environment data, etc.) unsupervised
[51] 2018 M vessel CPS OT (ship SCADA network) rule-based mathematical control theory CPS attacks
integrated with rules
[52] 2017 T vessel CPS OT (ship ICS and testbed dataset generation platform -
SCADA network) without any specific
monitoring method
[53] 2013 M distributed maritime network traffic unsupervised Maximum Entropy network
system/network Estimation layer attacks
Paper Type: M – proposes a monitoring method; A – proposes an architecture; T – proposes a testbed
When using Table IV for analysing the mapping of papers to taxonomy categories in Figure 2, we discovered four non-overlapping clusters of closely related papers. Each cluster contains five studies and the clusters are thus covering 20 papers out of 23. The clusters have been highlighted with different colours (red, blue, green, and orange) in Figure 2. If the majority of papers in some category belong to a single cluster, the category is highlighted with the colour of the respective cluster in Figure 2. Note that the “Distributed maritime system or network on land and sea” category in Figure 2 is highlighted with two colours, since it contains papers from two clusters which cover 50% and 30% of that category, respectively. Also, the “Unsupervised” category in Figure 2 is not highlighted with any colour, since it is the only category which contains papers from three clusters, and each cluster covers less than half of the papers in that category. The four clusters from Figure 2 are summarised below:
  • Studies on security monitoring for navigational systems [36, 40, 15, 45, 49] (red cluster in Figure 2) – papers which propose the analysis of application layer protocol data for detecting AIS and GPS spoofing attacks and anomalies in navigational systems.
  • Studies on security monitoring for CPSs [41, 43, 44, 46, 51] (blue cluster in Figure 2) – papers which propose the analysis of OT data for detecting attacks against vessel and canal CPSs.
  • Studies on security monitoring for distributed maritime systems and networks [34, 35, 37, 38, 53] (green cluster in Figure 2) – papers which propose the analysis of network traffic for detecting attacks against maritime IoT networks, maritime Software Defined Networks (SDNs), and navy networks.
  • Studies on security monitoring architectures and testbeds [14, 39, 47, 13, 48] (orange cluster in Figure 2) – the papers which have not proposed one particular security monitoring method like the studies from previous clusters, but rather a security monitoring architecture or testbed.
Detected clusters from Figure 2 reveal some associations between taxonomy categories which describe the nature of existing research. First, studies on security monitoring for navigational systems (red cluster) have focused on the analysis of relevant application layer protocols, and rule-based security monitoring approaches have been usually applied in the context of navigational systems. Furthermore, studies on security monitoring for distributed maritime systems and networks (green cluster) have focused on network traffic analysis, and such studies tend to employ supervised ML methods. Also, studies on security monitoring for CPSs (blue cluster) have proposed the analysis of OT data, and graph-based approaches have been employed only in this particular context. Unlike papers from other clusters, the studies on security monitoring architectures and testbeds (orange cluster) propose a range of different methods and their simultaneous use with the ability to process different types of input data and detect different attacks and anomalies.
It should be pointed out that three papers not covered by the above clusters are fairly close to them. Although in [50, 52] the authors have described an architecture and a testbed for processing OT data only, these papers are otherwise similar to the orange cluster in Figure 2. Also, although the paper [42] focuses on the analysis application layer data instead of network traffic, it proposes a security monitoring method for maritime IoT systems, and is thus closely linked to the green cluster in Figure 2. In the remainder of the paper, we will consider these three papers together with their closest clusters.
In order to distinguish papers that have proposed security monitoring methods from the papers that have described architectures and testbeds, we divided all papers by their type (see the “Paper Type” column in Table IV). Letter M denotes the papers that discuss specific security monitoring methods, letter A the papers which describe the architecture of the entire security monitoring system, and letter T the papers that discuss testbeds designed for security monitoring data collection, testing, and exercises. As can be seen from Table IV, 16 papers have focused on specific security monitoring methods [43, 34, 35, 49, 53, 15, 41, 44, 36, 37, 42, 40, 45, 51, 38], while 4 papers have described architectures [13, 48, 47, 50] and 3 papers have described testbeds [52, 14, 39], respectively. Most papers that concern architectures and testbeds [13, 52, 48, 47, 14, 39] have not proposed any novel security monitoring algorithms for analysing specific data (e.g., network traffic), but have rather employed various well-known industrial security monitoring technologies such as signature-based network IDS, antivirus, etc. which monitor a wide range of different data types (see the “Monitored Data” and “Monitoring Method” columns in Table IV). Therefore, we have set the “Method Type” to architecture or testbed for these papers in Table IV, because the focus of the paper lies on the entire architecture or testbed, not on one security monitoring method. The only exception is the work by Jia et al. [50] which describes the use of an Isolation Forest-based algorithm within the proposed architecture.
As can be seen from Table IV, security monitoring architectures received a significant amount of attention until 2019 (all 4 relevant papers were published in 2018–2019, constituting almost half of the 9 papers published before 2020). In contrast, during the more recent years after 2019, the research focus has shifted towards security monitoring methods and testbeds, with the majority of the papers (12 out of 14) discussing security monitoring methods. That finding illustrates the fact that during the earlier research period before 2020, scholars preferred to focus on the architecture of the entire security monitoring system, studying the unique challenges that arise in maritime environments. After the architectural issues and solutions for them were identified, the attention moved to particular security monitoring methods which constitute the building blocks of security monitoring system architectures.
The “Targeted Environment” column in Table IV indicates that most papers have proposed methods, architectures, and testbeds for addressing the needs of vessels or a distributed environment which involves vessels and land-based objects. In contrast, a small fraction of works target land-based environments only [46, 47].
As for the nature of monitored data (see the “Monitored Data” column in Table IV), we have used the term network traffic for studies which have proposed network monitoring methods without considering the application layer protocol in network packets. In the rest of the article, we will call such methods network traffic monitoring methods. Almost all the studies published until 2022 were not focusing on monitoring network traffic, with the work by Ptasinski et al. [53] being the only exception. For 2023, a new trend can be observed where a significant proportion of papers have described such methods [34, 35, 37, 38]. However, these studies have been limited to specific environments like maritime IoT networks and SDN systems. Also, according to Figure 2, the majority of papers that have proposed the analysis of application layer data have employed it for security monitoring of navigational systems, whereas the analysis of OT data is generally used in the context of CPS.
As for the types of security monitoring methods (see the “Method Type” column in Table IV), supervised and unsupervised ML approaches are the most common, while some papers have also described rule-based and graph-based approaches. As for supervised ML, deep learning-based approaches have been the most widely employed [34, 46, 37, 38], whereas no unsupervised ML method has seen more frequent use than the others. As Figure 2 illustrates, more than half of the papers which have proposed supervised ML approaches have applied them for maritime network security monitoring, whereas rule-based approaches tend to be applied for navigational systems. The only paper which has described a graph-based approach [44] has employed it for CPS security monitoring.
TABLE V: Overview of Research Papers on Security Monitoring Methods
Paper Evaluation Is Data Is Data Is Code Performance Computational Detected
Data Maritime Publicly Publicly Evaluation Cost Attacks and Anomalies
Specific Available Available Assessment
[34] NSL-KDD no yes no yes no network attacks against maritime IoT systems
[35] CICDDoS2019 no yes no yes yes DDoS attacks against maritime IoT systems
[36] MARSIM yes yes yes yes no GPS spoofing
[37] TON-IoT no yes no yes no network attacks against maritime IoT systems
[38] X-IIoTID no yes no yes no network attacks against
maritime SDN-based systems
[40] internal yes no no no no anomalies in NMEA messages
[15] internal yes no no yes yes GPS spoofing
[41] internal yes no no no no attacks against ship’s CPSs
[42] DS2OS no yes no yes no application layer attacks against
maritime IoT systems
[43] internal yes no no no no attacks against ship’s CPSs
[44] internal yes no no yes no attacks against ship’s CPSs
[45] internal yes no no no no anomalies in AIS messages
[46] internal yes no no no no attacks against canal lock CPS
[49] MarineTraffic AIS yes no no yes no AIS spoofing
[50] internal yes no no no no anomalies in the state and trajectory
of unmanned surface vehicle
[51] internal yes no no no no anomalies in ship’s CPSs
[53] internal yes no no no no network flooding attacks against navy networks
Table V provides an overview of 17 papers that propose security monitoring methods (i.e., the papers from Table IV with the “Method Type” column containing other values than architecture or testbed). Similarly to Table IV, Table V contains the “Detected Attacks and Anomalies” column, but provides more detailed information about the nature of attacks and anomalies. As this column indicates, some studies have proposed methods for detecting specific attack types. For example, whereas methods from [34, 37, 38] can detect network attacks of various types, algorithms from [35, 53] detect Distributed Denial of Service (DDoS) and network flooding attacks only. As another example, [40] proposes generic anomaly detection for NMEA messages, while [15, 36] focus on processing NMEA messages for the detection of GPS spoofing. Other methods described in papers from Table V include AIS anomaly detection [45], detection of AIS spoofing [49], detection of application layer attacks against maritime IoT systems [42], anomaly detection in the state and trajectory of unmanned surface vehicles [50], detection of attacks against canal lock CPS [46], and detection of attacks and anomalies in the CPSs of ships [43, 41, 44, 51].
Table V also provides some insights into the evaluation process of the proposed security monitoring methods. As can be seen from Table V, whenever maritime-specific data has been used for evaluating the method, the dataset is generally private (i.e., not publicly available). Also, whenever a public dataset has been used for evaluation, it is generally not maritime-specific. The only exception is the study by Spravil et al. [36], where the authors have created and released the MARSIM dataset.
Table V also reveals several other aspects of method evaluations – not all studies have published widely acknowledged performance metrics like precision, recall, and F1-score which provide a realistic picture of the method performance (for such studies, the “Performance Evaluation” column in Table V contains no). Also, most studies have not released the implementation of the method (see the “Is Code Publicly Available” column in Table V) and have not assessed the computational cost of the method (see the “Computational Cost Assessment” column in Table V).
The following sections provide a more detailed discussion of all papers from Table IV. For structuring the discussion, we have closely followed the aforementioned clustering of the selected research papers into four clusters based on the taxonomy presented in Figure 2. The papers on security monitoring architectures and testbeds (with the “Paper Type” column set to A or T in Table IV) are covered in Sections V-E and V-F, respectively (note that such papers correspond to the orange cluster in Figure 2). The discussion of the remaining papers on security monitoring methods (with the “Paper Type” column set to M in Table IV) has been divided into three subsections by the targeted environment (represented by red, blue, and green clusters in Figure 2) – the papers concerning security monitoring methods for distributed maritime systems and networks are covered in Section V-B, papers on security monitoring methods for navigational systems are covered in Section V-C, and papers on security monitoring methods for maritime CPSs are covered in Section V-D.
V-B Security monitoring methods for distributed maritime systems and networks
Gyamfi et al. [35] developed an adaptive incremental passive-aggressive ML (AI-PAML) method for a network attack detection system (NADS) in an IoT-based Maritime Transportation System (MTS) environment. A resource-efficient multi-access edge computing (MEC) setting was proposed to execute the system at network edges. To avoid data saturation problems of online learning models and facilitate the updating of the detection system, the authors provided an enhanced approximate linear dependence and advanced data updating technique to remove the irrelevant data. For efficient execution, the computation-hungry algorithms of NADS were deployed through MEC servers located on the ship or onshore. Markov Chains were then applied to facilitate the execution process, avoiding congestion on a MEC server, by predicting the next available MEC server with statistical reasoning.
The experimental setup for evaluating the work of Gyamfi et al. [35] was based on an actual IoT device network topology consisting of an RRH (Remote Radio Head for the IoT-MEC connection), MEC servers, and sensors-IoT device nodes in a vessel or at a port. The experimental results were obtained by using a PC running a layer 7 DoS simulator named DDOSIM. This was used to execute DDoS attacks against IoT devices with the aim being to achieve real-time responses from the developed NADS. The CICDDoS2019 dataset [^1] https://www.unb.ca/cic/datasets/ddos-2019.html was used to train the AI-PAML model. Although this dataset is quite recent and relevant for the IoT domain, it is not directly connected to the maritime domain. The authors also benchmarked the proposed model against the “Perceptron” and “Stochastic Gradient Descent” models regarding the training and prediction runtime and its latency was found to be better than the benchmark models after all models run against the DDOSIM attack.
Kumar et al. [37] proposed a monitoring system that uses Cyber Threat Intelligence (CTI) extracted with the deep learning-based method, Long Short-Term Memory-based Variational Autoencoder. The system applied the Bi-directional Gated Recurrent Unit (Bi-GRU) method for detecting attacks addressing IoT devices. The main drawback of the study is the fact that no maritime-specific dataset was used for the validation of the method, but rather the TON-IoT dataset [^2] https://research.unsw.edu.au/projects/toniot-datasets which contains network traffic of consumer IoT devices. Although the study mentioned that the proposed method is computationally more expensive than traditional ML algorithms and requires specialised hardware, the computational cost assessment was not conducted in the study.
Liu et al. [34] proposed a CNN-MLP based Intrusion Detection model for MTSs (Maritime Transportation Systems) which was trained via Federated Learning (FL), also called FedBatch by the authors. In the mentioned model, a Convolutional Neural Network (CNN) was utilised for data feature extractions, while Multilayer Perception (MLP) was used for attack classification to locally detect intrusion on each vessel. The study targeted IoT-based MTSs and endeavoured to address the limitations inherent in conventional learning algorithms, which consume an amount of computing resources that can not be provided in resource-constrained IoT environments. This study aims to protect data privacy by not sharing the training data obtained locally in vessels with the cloud. Instead, the models are induced locally in the remote devices, and their parameters are shared with the cloud. As the maritime environment may not enable all remote devices to submit their parameters in each learning round, the study utilises a batch federated aggregation approach. Thus, heterogeneous devices can upload their model parameters in a timely and synchronised way. The authors proposed a robust and efficient model for MTS security monitoring by training and testing their model on the NSL-KDD dataset. [^3] https://www.unb.ca/cic/datasets/nsl.html The model gained an overall performance accuracy of 88.1%, which surpassed the accuracy score of state-of-the-art CNN, i.e., 87.1%. Compared to other methods, the aforementioned method also achieved higher recall and almost the same precision on some attack types, i.e., Probe, R2L. Although much effort was devoted to researching FL-based intrusion detection in [34], one drawback of the study is the use of the NSL-KDD dataset, a general attack dataset, which lacks maritime domain specificity.
Ptasinski et. al [53] presented a Maximum Entropy Estimation based statistical ciphertext (CT) flow analysis mechanism to detect malicious flooding network attacks. The authors discussed the significant aspect of network security in the context of the Automated Digital Network System (ADNS) INC III, a critical network system of the US Navy for unclassified and secret information exchange. It interfaces between ships, shore-based Regional Network Operations and Security Centers (RNOSC) and Naval Computer and TeleCom Stations (NCTS). The authors specifically focused on protecting (Quality of Service) QoS in the CT domain against network flooding Denial of Service (DoS) attacks.
The mentioned research by Ptasinski et. al [53] provided an overview of different aspects of Plaintext (PT) and CT domain, i.e., QoS, CT Pcore, community of interest (COI) PT Enclaves, and showed the feasibility of flooding attacks. The authors defined normal traffic classes within the CT Pcore and also included the network plans, i.e., application plan, PT management plan, and security plan, within their proposed architecture. Moreover, the proposed approach detected anomalies in the network by using anomaly-free packets as training data and estimated their distribution as a baseline which represented normal traffic. Then the network traffic was observed and its distribution was compared with the baseline distribution. If the observed distribution was different from the baseline one, then an anomaly was reported. The authors performed testing by simulating packet flooding attacks on the Pcore QoS on the secret enclave group of traffic classes. Furthermore, the aforementioned study provided an attack response mechanism by limiting network traffic. However, in some cases, normal traffic flow was also unnecessarily limited.
In [38], Zainudin et al. introduced a decentralised trust aggregation solution which uses blockchain to provide a secure and trusted federated IDS structure. The solution offers a federated intrusion classification framework for SDN enabled marine traffic services. The scheme introduced a word wide available file system, with a network that uses blockchain for proof-of-authority (PoA). An IDS model that used FL-based collaboration was also proposed, together with the implemented, relatively simple intrusion detection and classification model. Their solution, leveraging CNN, surpassed other comparable implementations. The effectiveness of this model was evaluated using the X-IIoTID dataset, [^4] https://ieee-dataport.org/documents/x-iiotid-connectivity-and-device-agnostic-intrusion-dataset-industrial-internet-things specifically curated for IoT-related research.
Tiwari et al. [42] proposed a relatively simple but still resilient model to secure sensor networks against cyber attacks and adversarial activities. Their work focuses on the maritime sector’s IoT-based systems. The framework addressed the weaknesses in the network layer of the abovementioned infrastructure. The paper introduced the data collection, pre-processing, model training, method optimisation, testing and its results. According to the results, the Enhanced Light Gradient Boosting Machine (Light-GBM) technique delivered excellent results on low computational cost and bandwidth, which is essential in the huge and complex marine IoT environment, that is also limited from the communication bandwidth’s viewpoint. The experiments and the IoT attacks classification were conducted on the Distributed Smart Space Orchestration System (DS2OS) dataset, [^5] https://www.kaggle.com/datasets/francoisxa/ds2ostraffictraces an open-source but not maritime-related dataset.
Discussion. According to the analysis of 6 papers in this section, the current research on security monitoring methods for maritime networks has focused on specialised networks – IoT networks have been studied in four papers [34, 35, 37, 42], whereas two papers have targeted SDN [38] and military networks [53]. Moreover, the paper on monitoring SDN considered an IoT-specific dataset during evaluations, making IoT networks a dominant theme in relevant research. Therefore, the monitoring of other maritime network types than IoT has not received enough attention and constitutes an understudied research area.
As for the network monitoring methods, supervised ML methods have been the most frequently employed (proposed in five papers [34, 37, 38, 42, 35]), with deep learning-based approaches being the most prominent of them [34, 37, 38]. On the other hand, as argued in a recent domain overview study [17], the superiority of deep learning over traditional ML methods is not yet proven in the field of network intrusion detection, and deep learning involves a significantly higher computational cost which might not always justify its use for network monitoring. These considerations are even more pronounced for maritime systems which are often resource constrained. Therefore, the research on lightweight ML methods suitable for maritime networks is another understudied area.
Finally, the use of supervised ML for network security monitoring is known to suffer from the issue of concept drift [17] – since the surrounding environment changes over time, a supervised ML model will become obsolete, requiring retraining on new labeled datasets which are expensive to create. Existing works have not addressed this issue on datasets which cover longer time frames (e.g., several months), and studying ML methods for tackling concept drift (such as active learning) is a highly relevant future research direction. To this end, only the work by Gyamfi et al. [35] touched on this issue by providing a technique to keep only relevant data used for the dynamic updating of the learned model, although specific validation for this aspect of the work was not provided. Another interesting direction is further study of unsupervised ML methods, which do not need labeled data.
V-C Security monitoring methods for navigational systems
Kontopoulos et al. [49] presented a distributed architecture capable of detecting spoofing and falsification attacks in streams of AIS data in real time. Average speed is measured on the shortest path between positions obtained from two consecutive AIS broadcast messages. Then, this value constitutes the basis of the spoof detection approach implemented in this study. The distributed architecture was of the Master-Workers type, with the Master node distributing the AIS messages to Worker nodes for analysis in an efficient manner incorporating load balancing techniques.
The approach by Kontopoulos et al. was validated by generating a 43,912,236 AIS messages dataset from marinetraffic.com and performing simulations by attacking the dataset in various ways. The validation dataset has not been made public. The reduction of false positives as well as the extension to more attack types was mentioned to be part of the authors’ future work.
Amro et al. [40] proposed a method for analysing navigational NMEA-0183 messages. These messages communicate the data of the various onboard sensors. The objective of the method was to identify possible anomalies and their malicious causes. For this, the authors developed relevant anomaly detection algorithms. To facilitate the analysis the authors employed a maritime cyber security testing environment that they previously developed [54] which included the NMEA-Manipulator tool enabling the generation of the anomalies by attacking sensor data. The approach included a software package for creating synthetic datasets (which were not made publicly available) including normal and attack-related NMEA messages.
To develop the anomaly detection method, the authors first identified NMEA message types that affected navigation together with relevant message fields and their values. The authors also studied potential anomalies that can appear in NMEA messages and attack techniques which can trigger these anomalies, using the MITRE ATT&CK framework. [^6] https://attack.mitre.org/ Finally, the authors investigated different approaches for implementing attack and anomaly detection for NMEA messages. According to the authors, specification-based approach was best suited for that purpose, which involved the use of rules that described the normal behaviour of the system (i.e., any violation of these rules indicated an anomaly). Also, for some specific attacks that involved changing the arrival rate of NMEA messages to the system components, the authors recommended frequency-based anomaly detection that involved checking if the message rate remained within the expected boundaries. Although the study described the experiments for detecting a number of attack scenarios, precision and recall metrics were not reported in the paper (these metrics reflect the amount of false positives and false negatives).
The authors identified several limitations such as i) the possibility for false positives if the system is not well-configured, ii) the possibility of false negatives for sophisticated attacks that do not violate the rules of normal behaviour of the system, iii) not covering all the NMEA-0183 message types, iv) covering only NMEA-0183 and not other protocols, v) not considering all the possible anomalies. Finally, Amro et al. provided a discussion on possible deployment options of NMEA IDS on a vessel.
Iphar, Napoli and Ray [45] analysed the weaknesses of AIS in their study. Since these weaknesses allow for falsifying and spoofing AIS messages, the authors developed a rule-based system for anomaly detection from AIS messages. In their study, the authors proposed over 900 rules or integrity items for assessing the integrity of AIS messages. The first order assessment rules considered only one field from a single AIS message, while the second order rules analysed several fields from a single message. The third order rules considered fields from several AIS messages of the same type, whereas the fourth order rules analysed several AIS messages of all types. Based on the truth values of rules and external contextual information, flags were set that were further used for calculating ship-related risks and their levels.
For evaluating the proposed method, a Python-based prototype was created which utilised a database for storing AIS data, intermediate results from computations, and contextual data. The prototype was evaluated on about 24 million real-life AIS messages collected over 6 months. In the study [45], the authors provided a detailed description of risk evaluation for 4 cases involving anomalous AIS messages (the study considered 13 cases in total). In all 4 cases, the prototype raised one or more flags. However, the study did not evaluate the system on a larger number of cases, including the assessment of a false positive rate of the system (reflected by the precision metric in ML literature). Also, the developed prototype and the dataset were not made publicly available by the authors.
The study by Boudehenn et al. [15] focused on the detection of GPS spoofing. For that purpose, an anomaly detection approach was proposed which employed a One-Class Support Vector Machine (OC-SVM) and analysed NMEA messages that carried GPS information. The authors selected OC-SVM because it was easy to implement and did not require much computational resources, making it suitable for deployment on low-end hardware. During the experiments, the authors trained OC-SVM on data collected from a maritime simulator and also on real-life data collected on a small boat. For implementing GPS spoofing during the experiments, the authors utilised a device which generated a GPS signal carrying wrong GPS information.
According to the experiments, the OC-SVM based anomaly detection method featured a high recall and precision. To demonstrate the lightweight nature of the method and its low computational cost, the authors implemented the method on a Raspberry Pi based monitoring device that was suitable for deployment on real-life ships. Such evaluation of the actual computational cost of the proposed method and its suitability for real-life deployment sets this study apart from many similar studies which do not provide such information. The drawbacks of the study are the lack of publicly released experiment data and publicly available implementation of the method.
Similarly to the previous paper by Boudehenn et al. [15], the study by Spravil et al. [36] focused on the detection of GPS spoofing with the analysis of relevant NMEA messages. According to the authors, the analysis of NMEA messages allowed for the development of dedicated monitoring modules which could be easily integrated into the ship’s network. To demonstrate the viability of that approach, the authors publicly released the implementation of the MANA (MAritime Nmea-based Anomaly detection) framework. The framework combined unsupervised anomaly detection methods into an ensemble, and to identify suitable methods, the authors investigated various GPS spoofing methods, with a number of methods originating from other domains than maritime (e.g., aviation and mobile phone networks). As a result of the investigation, five methods were selected for the ensemble.
To assess the performance of the MANA framework, the authors created the MARSIM dataset which contained a large number of different GPS spoofing attacks and normal GPS data. Also, the dataset was released into the public domain, [^7] https://github.com/fkie-cad/mana making the study by Spravil et al. [36] one of the very few where a public maritime-specific dataset has been utilised for maritime cyber security monitoring research. According to the experiments, the ensemble of five methods featured a higher recall than each method (the ensemble regarded the anomaly as detected if any of the five methods provided a relevant indication). As one of the limitations of their approach, the authors identified the need for well-chosen thresholds for each method to avoid false positive alerts. Another limitation of the study is the lack of performance data describing the computational cost of the MANA framework.
Discussion. All five papers discussed in this section have focused on the analysis of application layer data in NMEA and AIS messages in order to detect AIS spoofing, GPS spoofing, and other navigational anomalies. As for the used methods, rule-based approaches have been suggested in three papers [49, 40, 45], whereas two papers have suggested unsupervised ML approaches [15, 36]. When employed for security monitoring, rule-based approaches are known to be very precise [17] and are thus used by many industrial security monitoring products (such as commonly used network IDS platforms like Suricata). However, creating a rule-based system requires domain experts and is a time-consuming process. One open research area is the use of rule mining algorithms in order to speed up the rule creation process.
Also, the current research has focused on rule-based and unsupervised ML-based approaches, whereas papers analysed in this section have not suggested the use of supervised ML for the detection of navigational attacks and anomalies. Therefore, studying the feasibility of supervised ML for that purpose is another open research area, as is dealing with advanced issues of supervised ML like concept drift. Finally, as mentioned in [17], rule-based and ML-based security monitoring systems are complementary, with both of them having unique advantages. Therefore, building and evaluating such hybrid systems for navigational security monitoring is another future research direction. Inductive logic programming [55] is a potential hybrid approach which could be applied to automate the rule learning process in a supervised ML manner.
V-D Security monitoring methods for maritime CPSs
Pelissero, Laso and Puentes [44] created and utilised graph-based models for the detection of attack propagation in CPSs. A graph model that represented the digital and physical system layers was proposed in the paper. It considered the interdependencies between CPSs by using the system variables. The model was applied in a case study that involved a small-scale system consisting of a ship with a fuel tank, valve, motor and propeller which were monitored through a PLC. One scenario of normal system usage and four attack scenarios including port scan, closing the PLC and man-in-the-middle attacks were simulated in the study. According to the authors, the method did not produce a false positive alert for the normal scenario, whereas one attack scenario out of four remained undetected (i.e., precision and recall were 100% and 75%, respectively). However, the evaluation of the method was limited to aforementioned five scenarios without testing the method on a larger dataset. The use of a significantly larger dataset would provide a better estimate about the long-term performance of the method when cyber attacks of a number of different types need to be handled. Furthermore, the applicability of the method would be questionable in large-scale systems as identification of the dependencies between system components is a non-trivial task that involves dedicated approaches [56].
Yoginath et al. [46] proposed a Recurrent Neural Network (RNN) based Digital Twin (DT) model to simulate an operational CPS of a Canal Lock. Their model performed real-time analysis to detect anomalies at the earliest, specifically anomalous changes in water levels within chambers through which ships pass. The study [46] used an actual PLC, sensors, valves and pumps to emulate the behaviour of the Canal Lock CPS physically. The authors identified all the possible scenarios, i.e., the ship’s up and down movements, performed by Canal Lock CPS and trained a set of RNN models using the emulation data as input. Anomaly detection was performed through a difference-based method (analysis of PLC and DT predicted values). The previously DT predicted value was compared with the current PLC read value and the deviation was captured, with the larger value being regarded as an indication of the attack. Furthermore, to evaluate the RNN model, the authors performed live validation of the system. However, any specific CPS attacks were not executed to test the accuracy of the DT concept for anomaly detection.
Xing, Cao and Chen [51] presented an anomaly detection method for the Ship Information System (SIS) based on risk data analysis. Since SIS is a typical SCADA system, the authors first elaborated on all the different operational and networking components of the SIS SCADA system and then they proceeded with presenting the data anomaly detection approach based on a cooperative state space control mathematical model, for the data propagated between the sensors, the distributed controller units and the actuators. A so-called Critical State Estimation (CSE) Algorithm was proposed based on the Industrial State Modelling Language (ISML). Detection rules were formulated as “condition implies action” with the condition being a boolean formula composed by various predicates indicating values that are expected by system components. The approach was validated by simulations on the ship dynamics under heading sensor signal attacks. The implementation of the controller was not made publicly available nor was the validation sensor data.
The study by Boudehenn et al. [43] focused on the use of Teager-Kaiser operator for time series analysis to detect anomalies in a ship’s propulsion system in the case of cyber attacks. During the experiments, the authors collected the following time series data – oil consumption, fuel consumption, propeller speed, and valve opening. The following attacks were conducted for evaluating the anomaly detection methods – port scanning with nmap tool, DoS attack with hping tool, and exploiting the vulnerability of the ship’s propulsion PLC for stopping and starting the PLC. Port scanning and DoS attacks did not influence the propulsion system, and only the attack against the PLC had a significant impact. According to the experiments, the Teager-Kaiser operator managed to identify sudden changes in the ship’s propulsion system time series data. However, the evaluation results were presented in the form of graphs only for the 5-minute time period which involves a single attack against the PLC. Therefore, it is difficult to assess what would be the detection rate for many attacks of different types against the propulsion system during a longer time frame (e.g., several days or weeks). Furthermore, the study did not include the analysis of false positive rates over a longer period of normal navigation without any cyber attacks (normal navigation can easily involve sudden changes in the propulsion system time series data, e.g., due to changes in engine speed by the crew). Although the study used a maritime specific dataset, neither the data nor the experiment code were released to the public domain by the authors. In addition, the study did not present any data on the computational cost of the anomaly detection method.
The study by Qiu et al. [41] proposed the use of a Particle filter-based anomaly detection method for time series data collected from a ship’s CPS to detect cyber attacks. In the study, the authors focused on data collected from the ship’s propulsion system – the temperature of the cylinder exhaust, cooling water pressure, and the speed of the fuel flow. For training the anomaly detection system, the authors used data collected from a new cargo ship during its first sailing week, assuming that the probability of cyber attacks during this week is very low. For evaluating the anomaly detection algorithm, data from the second half of the first sailing year was employed. For discussing the anomaly detection rate of the algorithm, the authors presented graphs for the 5,000-minute time frame (about 3.5 days) in the study. Although the graphs displayed some fluctuations in the time series data for the propulsion system, the study did not detail whether these sudden changes were the result of a cyber attack or rather a side effect of normal navigation. Also, the paper did not provide any experimental data that would involve real or simulated cyber attacks, measuring the precision and recall of the proposed anomaly detection method (traditional metrics for characterizing the performance of anomaly detection methods). In addition, the study did not provide any information about the computational cost of the method. Also, neither the experiment data nor the implementation of the anomaly detection method were released to the public domain.
Discussion. As for the research themes in five analysed papers, three papers focused on the vessel propulsion system [43, 41, 44], whereas the remaining two papers targeted vessel SCADA network [51] and canal lock CPS [46], respectively. Therefore, apart from the propulsion system, other parts of the vessel CPS have received little attention so far in the context of cyber security monitoring. Similarly, research on security monitoring for land-based maritime CPSs is scarce, and these research gaps are likely to be addressed by future works.
As for the types of security monitoring methods (see Table IV), no particular method types have been dominantly employed for security monitoring of maritime CPSs, and proposed methods have included supervised ML [41, 46], unsupervised ML [43], rule-based [51], and graph-based [44] methods. However, when analysing OT data from maritime CPSs, existing methods have generally not considered discriminating cyber attack related CPS malfunctions from CPS faults not caused by malicious cyber activity. Therefore, the root cause analysis for maritime CPS faults and anomalies is an unexplored research area which deserves closer attention.
Furthermore, process-aware attacks target CPSs to induce disruptions in physical processes [57]. Temporal characteristics of the processes observed on the sensors, actuators or control algorithms are utilised for the detection of malicious actions beyond the failures [57]. The reviewed papers do not address this monitoring approach comprehensively. Just focusing on the control functions in the physical space via process-aware approaches may not be enough to discriminate the usual system failures from malicious actions. Process-centric indicators should be correlated with other cyber-space indicators that reveal the prior actions of the attackers before compromising the control functions. To facilitate this research, it is necessary to generate datasets that include indicators about physical and cyber spaces.
Finally, in order to properly study the effects of cyber attacks on maritime CPS, a highly realistic lab environment is needed which would mimic a real-life CPS as closely as possible. Apart from one recent work [39] that will be discussed in Section V-F, the creation of realistic maritime CPS testbeds has been understudied, and can be regarded as a promising future research direction.
V-E Security monitoring system architectures
The study by Jacq et al. [13] proposed a maritime SOC architecture. The authors first provided an overview of the specific constraints of the maritime domain which complicate building a maritime SOC (for example, the limited network connectivity of the vessels with shore). The SOC architecture described by the authors took these constraints into account, consisting of vessel-based and shore-based parts. The proposed vessel-based functional blocks of the SOC were the following:
  • Network Connection Safety allowed to safely connect monitoring sensors to monitored systems (e.g., port mirroring for network monitoring sensors);
  • Network Probe Isolation was responsible for monitoring the network (e.g., a signature-based IDS), while being isolated from other sensors;
  • Local Preprocessor normalised and correlated events, to lessen the amount of data that needed to be sent to shore via network link via limited bandwidth;
  • Local Engine stored events on a ship for local alerting;
  • Ship Shore Manager acted both as a data cache and sender of data from a ship to shore, keeping the data in cache if there were issues with the network link, and being responsible for the link bandwidth management;
  • Cyber Situational Awareness Console provided a local simplified overview of the ship’s cyber state to the crew. The presence of the local onboard monitoring capability allowed for addressing scenarios where connectivity with the shore-based SOC was missing and the ship’s situation could not be assessed in a central land-based monitoring centre.
The proposed shore-based functional blocks of the SOC were the following. First, Ship Shore Manager and Bandwidth Manager were responsible for receiving the security data from ships and managing network connectivity with remote parties; Central Processor was responsible for filtering and normalising the received security data, and Data Store was a big data solution that stored the normalised security data. In addition to the aforementioned blocks, the shore-based SOC had other functional blocks that can also be found in traditional SOCs. For example, human operators could use specialised tools for big data analytics, searching the collected security data, creating visualizations, etc., to achieve maritime cyber situational awareness. In addition, SOC could be linked to external threat intelligence sources and incident response platforms. Finally, collected security data could be used for creating a so-called maritime-recognised cyber picture of the status of individual ships and the entire fleet. In their study [13], the authors mentioned that the previously described architecture has been thoroughly tested. Although no detailed information was provided about the nature of these tests and the performance of the architecture, the architecture was validated in a later independent study [14] by building a maritime exercise environment according to this architecture (see Section V-F).
Schauer et al. [47] proposed a security monitoring system architecture for ports which would integrate security data from two distinct domains – traditional IT systems (i.e., cyber assets) and physical assets of the port. The authors proposed to collect a wide variety of security events from IT systems, including firewall events, antivirus events from end user devices, and event logs of applications and servers. To monitor the physical assets of the port, the authors suggested data collection from physical access control systems, dedicated monitoring sensors which can create security events, etc. According to the authors, security events from cyber and physical assets should be collected into the hybrid situational awareness framework which can analyse events from both domains.
For real-time analysis of these events, two major components were proposed for the situational awareness framework – Event Correlation Engine (ECE) and Threat Propagation Engine (TPE). According to a widely used definition [58], event correlation is a real-time process that involves assigning a new meaning to event groups which occur within predefined time windows. For implementing ECE, the DROOLS engine was proposed in [47]. As the authors stated, the use of ECE allows the detection of complex security incidents which manifest themselves through several events from different domains. The purpose of the TPE was to establish the impact of high-priority security alarms via graph-based analysis, where the nodes of the graph represented cyber and physical assets and the edges of the graph reflected the dependencies between assets.
The approach by Schauer et al. [47] was sound and novel considering the time of its publication, integrating cyber and physical situational awareness methods for port infrastructures (including communication with ships). On the other hand, the authors did not provide a deeper analysis of the possible cyber attacks and how they could be detected in the maritime infrastructure domain. Also, the approach was heavily human expert-based with regards to the creation of the pattern-based rules for the event correlation engine and also to raising alarms for triggering the operation of the threat propagation engine. This may contribute to biased decision making. Furthermore, the authors proposed the DROOLS engine for event correlation purposes without comparing it with other solutions (e.g., [59, 60]) to justify the selection of DROOLS. Last but not least, the approach was exemplified through two attack scenarios but no simulation results were provided and the starting of the attacks was assumed to succeed rather easily.
The paper by Jia et al. [50] discussed a big data architecture to collect OT data from Unmanned Surface Vehicles (USVs) for various purposes like business management, disaster prevention, anomaly detection, etc. The authors proposed the collection of the following data from USVs – ship’s performance and navigational data (e.g., speed, position, level of fuel, etc.), data about the environment (e.g., wind, temperature, etc.), data from sensing devices that provided information about the objects around the ship (e.g., data from the radar), and data that was collected from other sources than USV itself (e.g., industrial data about the USV collected from the manufacturer). For storing the collected data, a cloud-based solution was proposed that was PostgreSQL based. According to the authors, visualisation capabilities were an important aspect of the big data architecture, and the authors suggested the use of the LinDA toolkit for processing the data in the PostgreSQL database.
As an example use case for the big data architecture, Jia et al. described an anomaly detection scenario for identifying abnormal combinations of the sailing angle and the USV speed. For anomaly detection purposes, Isolation Forest was proposed. However, the authors did not provide a performance evaluation of the Isolation Forest based anomaly detection method on some datasets. Also, the code and the data used during the experiments were not made publicly available by the authors.
In [48] Möller et al. discussed an agent-based intelligent maritime traffic management system to support a coastal Vessel Traffic Service (VTS) with a clear situational picture, traffic and environmental information. For that purpose, the authors proposed to introduce a Multi-Agent System (MAS) that extended to vessels to exchange relevant data (e.g., navigational and safety information) with the central VTS. Following the concept of connected cars, the authors also introduced the concept of connected ships which were continuously in contact with each other. That facilitated the exchange of important navigational information not only through VTS but directly between ships.
The authors also described how to augment the proposed distributed MAS with a security monitoring functionality. First, the authors detailed the weaknesses of the wireless communication networks used by MAS, and mentioned a mitigation opportunity in the form of anomaly detection for data flows in the network. However, the authors did not provide a more detailed discussion on the anomaly detection algorithm suitable for that goal. Möller et al. also suggested the use of neural networks for the detection of cyber attacks. However, the discussion of the proposed method remained brief and no implementation information was provided. For example, although the study mentioned training neural networks on probability-based features, the nature of the input data was not detailed. Furthermore, the paper did not report any experiments with the proposed method together with attack detection performance results. For these reasons, it is difficult to assess if the suggested neural network based cyber attack detection method would be a viable choice for maritime systems.
Discussion. From the four analysed papers, two papers have proposed generic security monitoring architectures, with [13] discussing a maritime SOC and [47] a situational awareness system for a port infrastructure which is similar to SOC. The remaining two papers have proposed security monitoring architectures for specific purposes (a big data architecture for monitoring USVs [50] and a VTS-centric security monitoring architecture [48]).
Although the creation of SOCs is a widely studied cyber security topic (see [16] for a recent domain overview), most works in this area lack detailed recommendations and guidelines for implementing a SOC in a scalable and resource-efficient way (for example, recommendations on software solutions to use and guidelines for configuring them) [61]. Furthermore, only a few works have evaluated the SOC architectures and provided relevant performance data (e.g., for resource consumption and event processing rate of system components). Since maritime environments are often resource-constrained, these topics are also relevant in the maritime context and constitute valuable future research directions.
According to recent studies [62, 16], SOC analysts are often overwhelmed by security alerts of low importance which leads to analyst burnout and alert fatigue. For supporting the human analysts, alert prioritisation has been identified as an important topic for SOC environments [16], and several ML approaches have been recently proposed for that purpose [63, 64]. Since the employment of alert prioritisation for reducing human alert fatigue has not been studied in the context of maritime SOC, it can be regarded as one of the open issues for future work.
V-F Security monitoring testbeds
The short paper authored by Becmeur et al. [52] aimed to present a platform to generate data and scenario traces for the evaluation of algorithms for intrusion detection. The introduced solution was based on a ship’s two critical subsystems: 1) The propulsion and engine control in which a computer fan simulated the ship’s propeller and the fan’s actual RPM represented propulsion. 2) The navigation subsystem with the rudder control to simulate the ship’s direction changes. Each subsystem had an independent controller which used common industrial communication protocols, like Modbus, DNP3, and S7. The heart of the setup was the central controller built on a Raspberry Pi. The simulated sensor – that provided the measurement of the engine parts’ temperature – was made of an open-source electronics platform called Arduino. The authors provided an overview of the project, elucidating its overarching objectives and outcomes, but the overview lacked the required depth in elucidating the procedural intricacies that are essential for comprehensively understanding and, crucially, reproducing the results.
The paper by Raimondi et al. [14] described a testbed for conducting cyber exercises to train maritime SOC teams. The testbed followed the maritime SOC architecture proposed by Jacq et al. [13] (see the previous section for a detailed discussion) and was implemented with Linux containerisation techniques. The testbed employed a ship simulator for simulating the ship at sea and a custom Python script for transmitting data received from the simulator as NMEA messages to other ship-related testbed components. For monitoring the ship network, the authors proposed Suricata IDS which employed Lua scripting for parsing NMEA messages. For the cyber situational awareness console on the ship, the authors used Splunk which also forwarded data to a shore-based central SIEM server (another instance of Splunk). To illustrate the training process of maritime SOC operators, the authors described an example cyber exercise which involved the injection of false NMEA messages into the ship network that interfered with the gyrocompass. The task of the trainees was to detect this attack with the help of Splunk query language.
The paper by Sicard et al. [39] described a testbed for experimenting with the Industrial Control System (ICS) of a warship. To achieve a high degree of realism, the authors utilised the physical devices of real ships as much as possible. The testbed implemented the following four areas of the warship – propulsion system (engine and propellers), direction (rudders), energy (fuel for the engine and fuel pump), and artillery (76mm main gun and motors for moving the gun turret). For these four areas, a physical implementation contained a bow and rear of the ship with a physical gun turret, rudders, propellers, physical devices of the bridge, etc. According to the authors, the use of virtualised solutions would have decreased the realism of the testbed, since they would have not allowed for the full imitation of the ship’s real-life environment. For low-level monitoring of Modbus RTU fieldbus traffic in the ICS of the ship, the authors proposed the use of Zeek IDS with specialised traffic capturing hardware and Zeek protocol parser for the Fieldbus protocol. For regular network monitoring, commercial IDS sensors were suggested by the authors. Although the study mentioned that the testbed could be utilised for generating maritime cyber security datasets and described four possible cyber attack scenarios, such datasets were not published by the study.
Discussion. From the three analysed papers, two papers have proposed testbeds for experimenting with vessel ICS and SCADA networks [52, 39], whereas [14] focuses on a SOC testbed for training maritime SOC analysts. Since [52] is a short paper, the discussion of the proposed testbed remains fairly brief, but in the remaining two papers [14, 39] detailed testbed descriptions are provided, with both testbeds being highly realistic environments for conducting maritime cyber security experiments and trainings.
However, it should be considered that the aforementioned studies represent specific environments (a warship’s ICS and a maritime SOC). Since for other maritime systems there are no similar studies, the creation of highly realistic testbeds for these systems constitutes an open research area. Also, realistic testbeds allow for generating high-quality datasets for maritime cyber security experiments. As discussed in Section V-A, there are very few publicly available maritime-specific datasets (see Table V), and the creation of such datasets is another important future research direction.
VI Discussion
In order to answer RQ3 (see Section III-A), this section presents the discussion of our main findings about the identified research gaps. The main findings are based on the analysis from the previous section and are presented below.
Finding 1: Lack of evaluations on maritime-specific datasets which are publicly available. Numerous publications have evaluated the proposed security monitoring methods on datasets that are not publicly accessible (see Table V), so the experiments in question are characterised by an inherent lack of reproducibility, thereby precluding the possibility of result verification. Although some studies have employed publicly available datasets, these publicly available datasets are not maritime specific but rather generic intrusion detection datasets like NSL-KDD, CICDDoS2019, TON-IoT, etc. (see Table V). Therefore, the validity of these research results remains questionable, because the datasets do not represent the actual maritime systems closely enough. The only exception is the study by Spravil et al. [36] which employed the publicly available maritime-specific MARSIM dataset.
Finding 2: Lack of data generation about system failures. According to recent domain overview papers [65, 66, 67], research datasets used for developing and testing security monitoring solutions in IT networks [65], ICSs [66], and CPSs [67] include attack and normal system activities. Thus, the developed security monitoring approaches discriminate attacks from normal activities or identify anomalies originating from malicious actions. We have observed a similar approach in the datasets related to the maritime domain. However, in practice, system failures arising from unintentional causes (e.g., software or hardware errors, environmental condition changes) can be mistakenly detected as anomalies resulting from cyber attacks. Thus, the root causes of the incidents are not usually identified accurately and timely, causing delays and discrepancies in incident handling and recovery operations. As pointed out in [67], it is important to include activities representing all failure modes of the target system in the datasets. For example, recent studies in the energy domain aim to discriminate cyber attack related anomalies from non-malicious ones [68, 69, 70]. Therefore, we contemplate that the generation of datasets that include system failures in addition to normal and attack cases would facilitate the research in developing solutions for more granular and informative security monitoring functions.
Finding 3: Lack of publicly available experiment code and prototype implementations. Most studies have not released the implementations of proposed methods and the experiment code (see Table V), with the only exception being the study by Spravil et al. [36]. This shortcoming is another serious obstacle to achieving reproducibility, and it also complicates the evaluation of the proposed methods on different datasets and in live environments.
Finding 4: Lack of proper performance evaluations with appropriate metrics. Several studies have failed to properly evaluate the performance of the proposed security monitoring methods with proper metrics (e.g., [41, 43, 45, 50, 46, 53]). However, in the field of security monitoring and ML widely acknowledged performance metrics exist such as precision, recall, and F1-score [17]. The failure to use such widely used metrics makes it difficult to assess what is the attack detection rate and false positive rate of the proposed methods, and how the proposed methods compare to other approaches.
Finding 5: Lack of proper computational cost evaluations. Only a few studies ([15, 35]) have assessed the computational cost (e.g., CPU time and memory consumption) of the proposed security monitoring methods. However, the computational cost is an important consideration, since vessel networks can usually not accommodate specialised computing platforms for expensive calculations. For example, some recently proposed security monitoring methods are based on deep learning [34, 46, 37, 38] which is known to be computationally expensive [17], and without the computation cost assessment the applicability of these methods in maritime environments remains questionable.
Finding 6: Lack of detailed assessments of cyber attack impacts on maritime CPSs. Some papers (e.g., [46, 41]) which have proposed methods for detecting cyber attacks against maritime CPS have not conducted cyber attacks in a relevant lab environment. However, the lack of such cyber attack experiments prevents to assess what is the real impact of these attacks on CPS, and what are the best avenues for detecting these attacks. Furthermore, the assessment of cyber attack impacts and best attack detection approaches requires a lab environment which mimics CPS as closely as possible. However, apart from one recent study [39], the methods for building highly realistic maritime CPS labs and testbeds have not received enough attention in the relevant literature.
Finding 7: Lack of detailed evaluation of the proposed security monitoring system architectures. Security incident handling processes usually require collecting, aggregating and correlating security events from various sources to do more relevant prioritisation, accurate incident categorisation and impact assessment. Although the idea of using maritime SOC and aggregating diverse security data from many sources has been proposed in several past studies [13, 47, 48, 50], existing works have not published detailed evaluation data for the proposed architectures. However, some similar studies that are not maritime specific have included such evaluations (e.g., see [61]), and works from maritime domain should follow the same approach.
Finding 8: Security monitoring methods for some maritime systems have been understudied. As discussed in Section V, security monitoring methods for distributed maritime networks have been largely focusing on IoT networks, while other network types have received less attention. Also, most network monitoring research has utilised supervised ML methods, and studying the feasibility of other methods is a valuable future research direction. As for maritime CPSs, systems apart from vessel propulsion have been less studied. Also, process-aware maritime CPS monitoring which correlates indicators from the physical and cyber space remains an open research topic. As for security monitoring methods for navigational systems, the use of supervised ML, rule mining, and inductive logic programming are other possible future research directions.
Finding 9: Advanced ML issues have been understudied in the maritime context. The lack of domain-specific datasets in the maritime field is a major obstacle to understanding the performance of ML methods in solving maritime security monitoring problems. However, once datasets are created, various interesting problems related to the application of ML methods can be addressed. Handling the concept drifts that may occur in the feature space, learning from small labelled datasets or efficient utilisation of experts for labelling are some research directions that can be tackled in this domain. The application of large language models for a better understanding of the detection rules or monitoring results [71] and receiving suggestions from these models about the course of actions can be studied in the maritime context. Some of the security monitoring tasks can be deployed into the resource-constraint devices in maritime systems for making real-time decisions. Thus, another potential research dimension would address running and optimising learning models on such devices (i.e., TinyML applications [72]).
Finding 10: Human aspects of security monitoring systems have been understudied in the maritime context. Although the ultimate goal of cyber security monitoring would be the full automatisation of attack detection and response mechanisms, it is highly expected that human experts will take a critical part in incident handling processes in any case. Thus, interdisciplinary and holistic approaches involving human and technical aspects of the problem domain are highly needed. Training and awareness programs for security analysts and maritime operators addressing design issues for the minimisation of human errors (e.g., optimising user interfaces of monitoring tools), enhancing communication between incident handling teams or selecting the right course of action under stressed conditions are some sample topics that require further research in this domain. Another potential research topic is the use of ML tools for alert prioritisation that helps to prevent alert fatigue among security analysts (a recent study among maritime cyber security experts has confirmed the importance of this topic [73]). Apart from some works which have discussed the expected competence of maritime security analysts and their training process [14], and the importance of cyber security competence among vessel crews [74], existing studies have generally not touched on human aspects of maritime security monitoring.
VII Conclusion, limitations and future work
In our research, we conducted a systematic literature review in which we surveyed four digital libraries, executed a manual search, and snowballed. After a careful selection process described in Section III, we identified 23 papers in the field of cyber security monitoring in maritime. In Section V, we provided a comprehensive overview of these publications, highlighting their strengths and shortcomings. We shared the bibliometric analysis of the papers in Section IV, and we discussed the identified research gaps in Section VI.
Our developed taxonomy (outlined in Section V-A) is a valuable tool for researchers and authors alike. It provides a structured framework for analysing cyber security monitoring-related literature, affording insights into the strengths and limitations of existing publications while concurrently delineating potential areas for future research. This taxonomy’s utility lies in its capacity to comprehensively encapsulate multifarious facets of the literature and its simple representation, rendering the analytical outcomes self-explanatory and enhancing accessibility for a diverse readership.
As we discussed in Section VI, we identified several shortcomings and limitations of the currently available literature. The present study yields the following findings.
First, we found several dataset related issues (Findings 1-2). Maritime-specific publicly available datasets should be used for experiments to ensure the reproducibility and the validity of research results for maritime environments. Facilitating the accessibility of publicly available maritime-related datasets would serve as a valuable resource for researchers, aiding and enhancing their investigatory endeavours in this domain. Also, the datasets should include system failure data not associated with cyber attacks, since that would allow us to evaluate how well the security monitoring methods can distinguish the effects of cyber attacks from other system failures.
Second, we identified several evaluation related issues (Findings 3-7). To improve the quality of evaluations, appropriate metrics should be used for assessing the performance of algorithms. Also, it is important to assess the computational cost of the proposed methods. Evaluations should be realistic, involving real cyber attacks that are conducted in a lab environment with a high degree of realism. For the sake of reproducibility, it is essential to share the experiment code and prototype implementations. Similarly to security monitoring methods, detailed evaluations are needed for security monitoring system architectures.
Third, we identified several previously unexplored research areas of maritime security monitoring (Findings 8-10). For example, the monitoring methods for maritime IoT networks and vessel propulsion systems have been studied more in the current literature, whereas other types of maritime distributed networks and maritime CPSs are potential targets of future research. Similarly, process-aware model-based methodologies, emphasising the correlation between physical and cyberspace indicators for enhanced system malfunction and cyber attack distinction, could be leveraged in the maritime domain. Furthermore, several advanced ML issues like concept drift and learning models for resource-constrained environments deserve more attention in the field of maritime security monitoring. While the training of personnel constitutes a fundamental aspect of cyber security, it is noteworthy that, upon thorough investigation, almost no scholarly literature specifically addressing this subject matter could be identified in the extant literature.
According to our assessment, these results fill the identified research gap and help the community in their research.
Although a comprehensive examination was undertaken through an exhaustive review of relevant literature, and the applied methodology facilitated the scrutiny of over 1000 sources, it is imperative to acknowledge the inherent limitation in ensuring the absence of any overlooked publications.
As for future work, we plan to study the methods of creating a realistic maritime lab, using our experience in the field [75]. Also, we plan to use this lab environment for the creation and release of maritime-specific datasets to facilitate security monitoring research in the maritime domain. Finally, our plans include studying advanced ML-based security monitoring methods on these datasets.
Acknowledgment
This research was funded by the EU Horizon2020 project MariCybERA (agreement No 952360).